OpenStudy (anonymous):

if du/dx=2x+6 what does du equal? this is using substitution for solving an indefinite integral.

OpenStudy (anonymous):

i guess \[du=(2x+6)dx\] what is the actual problem?

OpenStudy (anonymous):

du = (2x + 6)dx You can now integrate both sides to solve for u

OpenStudy (anonymous):

\[u = x^2 + 6x + C\]

OpenStudy (anonymous):

\[\int\limits_{?}^{?} x+3/(x^2+6x)^2 dx\]

OpenStudy (anonymous):

oh ok \[u=x^2+6x\] \[du= (2x+6)dx\] \[\frac{1}{2}du=(x+3)dx\] and then you get \[\frac{1}{2}\int \frac{1}{u}du\]

OpenStudy (anonymous):

a set up for a u - sub. then anti derivative is \[\frac{1}{2}\ln(u)\] , replace u by \[x^2+6x\] and you are done

OpenStudy (anonymous):

omg thank you sooooo much!!!